OpenStudy (anonymous):

find the area of a square if perimeter=128 .......

OpenStudy (anonymous):

find the side length of the square

OpenStudy (anonymous):

X+X+X+X=4X 4X=128 X=128/4=32 AREA=X.X=32*32=1024